# Roundtable Policy: Improving Scientific Reasoning and Narratives through Confidence-Weighted Consensus of LLMs

### v1

@@ 3.1 Performance Improvement (RQ1)
SingleTask |
GPT-4o | 70.995.01 | 74.734.68 | 82.753.76 | 75.495.23 | 74.815.03 | 79.634.31 | 72.155.19 | 70.535.25 | 79.644.19 |
| GPT-o3 | 84.073.98 | 84.253.82 | 86.463.42 | 85.923.95 | 86.703.56 | 85.843.56 | 83.904.15 | 81.524.25 | 84.863.69 | |
| GPT-4.1 | 76.964.62 | 81.374.16 | 86.193.51 | 81.654.67 | 83.194.19 | 84.753.85 | 78.814.74 | 77.914.67 | 83.903.88 | |
| GPT-4.1n | 77.394.55 | 80.944.15 | 85.593.53 | 80.484.69 | 82.304.19 | 83.973.85 | 77.974.74 | 77.224.65 | 82.943.90 | |
| GPT-o3-m | 82.654.16 | 84.783.83 | 87.263.36 | 83.154.35 | 86.863.53 | 86.443.53 | 81.504.40 | 81.824.23 | 85.153.67 | |
| GPT-4o-m | 72.784.94 | 74.644.70 | 82.193.80 | 75.445.16 | 76.184.85 | 79.814.23 | 72.785.18 | 70.695.24 | 78.824.25 | |
| Grok-3 | 78.505.10 | 79.104.50 | 85.503.64 | 80.964.75 | 81.264.47 | 83.583.99 | 78.814.78 | 76.344.93 | 83.044.01 | |
| Claude-3.7S | 82.474.22 | 84.733.92 | 88.173.28 | 85.524.20 | 86.653.79 | 86.893.57 | 84.534.22 | 82.134.38 | 86.413.59 | |
| DeepSeek-chat | 84.054.20 | 83.744.12 | 88.283.19 | 84.214.31 | 84.584.13 | 86.083.58 | 83.684.23 | 81.564.48 | 86.503.46 | |
| RP | 90.253.95 | 89.953.78 | 91.043.09 | 90.393.91 | 91.283.35 | 91.183.20 | 90.153.95 | 89.863.95 | 90.993.26 | |
To answer RQ1, we compare Roundtable Policy (RT) against individual LLM baselines across both MultiTask and SingleTask settings. Table 1 summarizes overall performance, while Figures 2(a)–2(b) visualize relative gains. This correspondence between player ID and model follows Table 3. The correspondence of subtasks of MultiTask and task IDs follows Table 4.
Quantitative Analysis.
In the MultiTask regime, RT delivers a 13.01% average lift, outperforming 8 out of 9 single-model baselines. The only exception is Task 9 (Bio.), where Qwen-plus () and GPT-4o () slightly exceed RT (). Beyond accuracy, RT reduces uncertainty; for instance, on Task 7 (Math) the confidence interval contracts to compared with baselines’ –. Figure 2(a) further shows per-player improvements. Gains are broadly positive on Tasks 1–8, with especially large margins on reasoning-intensive subtasks—for example, +90.5% over Gemini-2.0 on Task 7. Smaller improvements occur when a single model already aligns closely with ground truth, while Task 9 shows marginal or negative differences—consistent with its per-task averages. Appendix §D.1 confirms that weaker players (e.g., Gemini-2.0) enjoy the largest lifts, while stronger models exhibit limited headroom.
In the SingleTask regime, which evaluates structured scientific proposal writing, RT achieves the best performance in all nine rubric cells, explaining the observed 11.04% average lift. Each proposal comprises three sections—Background, Methodology, and Impact—scored across Creativity, Scientific Rigor, and Logical Coherence. Scores cluster tightly around 90–91% with narrow confidence intervals (typically –). Gains concentrate in the Impact section, particularly Scientific Rigor, as shown in Figure 2(b). Appendix §D.2 reveals positive medians across most players, with right-skewed tails indicating that even when consensus does not outperform the strongest model, it still elevates weaker drafts significantly.
Qualitative Analysis.
These results illustrate how RT acts like a committee process rather than a flat average. In MultiTask, the confidence-weight table down-weights unreliable contributions and amplifies complementary strengths, thereby “rescuing” weaker players without diluting strong ones. This mirrors how, in a committee, flawed proposals can still contribute useful fragments when evaluated critically. The large gains on reasoning-heavy tasks (e.g., Tasks 6–7) highlight how diverse error patterns across models can be combined into substantially stronger reasoning chains.
In SingleTask writing, weaker players show the most pronounced gains in Impact and Methodology. For instance, GPT-4o exhibits long right tails in Impact–Scientific Rigor and Impact–Logical Coherence, while GPT-o3 shows consistently positive medians with bounded variance. Even strong baselines such as DS-chat rarely degrade under RT, suggesting that consensus seldom harms high-quality drafts but frequently elevates weaker ones. Mechanistically, the confidence-weight table enforces rubric-aware, non-redundant fusion: systematically weak contributors are down-weighted, while demonstrated strengths are amplified. The pattern resembles how peer reviewers specialize—some focusing on methodology, others on impact—and the editor synthesizes these perspectives into a well-rounded final decision.
3.2 Long-Run Stability and Robustness (RQ2)
Setting Player ID LLM # Mean Gain (%) Std (%) p-value Significance MultiTask 001 GPT-4o 36000 6.476865 77.280369 ✓ 002 Grok-2 36000 5.254293 78.231667 ✓ 003 Claude-3.5Haiku 36000 17.251754 101.839946 ✓ 004 GPT-4.5-preview 36000 18.552776 94.463211 ✓ 005 Deepseek-Chat 36000 16.048551 96.935292 ✓ 006 Claude-3.7Sonnet 36000 13.732040 121.321577 ✓ 007 Llama-3.3 36000 24.988341 123.678543 ✓ 008 Qwen-plus 36000 12.586367 91.751670 ✓ 009 Gemini-2.0 36000 46.474125 145.071773 ✓ SingleTask 001 GPT-4o 36000 27.087781 40.303119 ✓ 002 GPT-o3 36000 13.316247 148.785675 ✓ 003 GPT-4.1 36000 15.107889 73.600957 ✓ 004 GPT-4.1-nano 36000 15.805949 28.188351 ✓ 005 GPT-o3-mini 36000 9.687830 21.760829 ✓ 006 GPT-4o-mini 36000 27.206693 42.707340 ✓ 007 Grok-3 36000 16.485277 29.024328 ✓ 008 Claude-3.7Sonnet 36000 7.761157 16.993190 ✓ 009 Deepseek-Chat 36000 7.165863 7.474061 ✓
To answer RQ2, we test whether RT’s gains remain statistically reliable over long-run evaluations. Using the Wilcoxon signed-rank test (§A.2) on round-level paired differences, RT consistently outperforms all baselines. As shown in Table 2, across all 18 models RT achieves significant improvements at , with -values typically below .
Quantitative Analysis.
In the MultiTask regime, RT consistently outperforms baselines with mean percentage gains ranging from +5.3% (Grok-2) to +46.5% (Gemini-2.0). Even strong players such as GPT-4o (+6.5%) and Claude-3.7Sonnet (+13.7%) see reliable gains, highlighting that consensus remains beneficial even when individual agents are already competitive. The large improvement over Gemini-2.0 illustrates how RT effectively leverages complementary strengths to elevate weaker contributors, rather than letting them dilute overall performance.
In the SingleTask regime, RT delivers significant lifts across all nine baselines, with mean improvements from +7.2% (Deepseek-Chat) to +27.2% (GPT-4o-mini). Notably, gains are especially pronounced for smaller or lighter-weight models (e.g., GPT-4o-mini, GPT-4o), suggesting that RT acts as a stabilizing mechanism that brings their drafts closer to the quality of stronger models. Even against advanced systems like GPT-o3 and Claude-3.7Sonnet, improvements remain consistently positive, confirming robustness across a wide range of capability levels.
Qualitative Analysis.
Beyond statistical significance, the long-run results highlight important qualitative patterns. In the MultiTask regime, RT’s robustness stems from its ability to repeatedly exploit complementary error profiles: weaker players such as Gemini-2.0 or Llama-3.3 often fail in different ways, but when their contributions are weighted through the confidence table, the ensemble consistently extracts correct elements while suppressing noise. This is reflected in the large variance reported in Table 2—although individual rounds fluctuate, the aggregated effect over thousands of rounds converges to stable gains. In committee terms, even if certain members are unreliable in a given discussion, their occasional correct insights are preserved while systematic mistakes are filtered out, ensuring dependable group performance.
In the SingleTask regime, stability is most visible at the rubric level. Analytical dimensions such as Impact–Scientific Rigor show the largest sustained gains, whereas stylistic traits (e.g., Creativity or Logical Coherence) improve more modestly. This asymmetry suggests that RT is especially effective in consolidating rigorous, reasoning-heavy contributions—akin to how editors place greater weight on technical soundness than on stylistic flair when synthesizing reviewer reports. Importantly, the consistency of improvements across tens of thousands of rounds demonstrates that RT does not rely on one-off synergies, but instead behaves like a well-tuned review board: repeatedly integrating diverse perspectives into a balanced and accountable consensus.
3.3 Bias and Consistency of AI Graders (RQ3)
To address RQ3, we examine grader bias and inter-grader consistency through two complementary lenses: (i) the distribution of ranks assigned to RT by different graders, and (ii) pairwise agreement measured by Kendall’s Tau (see §A.3). Aggregated results are shown in Figure 3, with task- and dimension-level breakdowns provided in Figures 14 and 13 in §D.3.
Quantitative Analysis.
Figure 3 highlights a clear distinction between MultiTask and SingleTask regimes. In MultiTask, the aggregated rank distribution is steep and concentrated in the top three positions, with RT frequently ranked first. Correspondingly, Kendall’s Tau values are consistently high (–), reflecting strong alignment among graders. Task-level inspection (Figures 14, 13) confirms this pattern: Tasks 001–006 exhibit overlapping distributions, Tasks 007–008 achieve near-perfect consensus (), while Task 009 shows the weakest agreement (–) as some graders disperse RT’s rank across mid-range positions.
In SingleTask, however, the aggregated rank distribution is flatter and more dispersed, and Kendall’s Tau values drop markedly (– on average). Creativity-related dimensions (e.g., Background–Creativity, Methodology–Creativity) place RT near the top but with substantial grader variability. Logical coherence dimensions show the lowest alignment, with Tau values as low as –, indicating inconsistent judgments. By contrast, Impact–Scientific Rigor yields the strongest agreement, with RT concentrated near rank 1 and Tau values of –. Together, these results demonstrate that graders achieve high consensus in objective, structured tasks but diverge significantly when assessing subjective, open-ended proposals.
Qualitative Analysis.
These results reveal that bias is minimal in structured MultiTask settings, where evaluation criteria are well defined, but more pronounced in open-ended SingleTask writing, where subjective interpretation shapes grading. This mirrors human peer review: reviewers evaluating technical correctness tend to agree, while judgments about creativity or style diverge. Crucially, the AI Committee mitigates such variability by aggregating across graders, reducing the impact of individual biases. In practice, this means RT functions like an editorial board: even when reviewers disagree, their collective judgment yields a more balanced and reliable outcome.
4 Ablation Study
To disentangle the contributions of our design choices and validate their scientific relevance, we conduct two ablation studies. These analyses illustrate how Roundtable Policy strengthens structured reasoning and improves narrative quality by highlighting what is gained—and what is lost—when key components are removed.
Policy Ablation.
We first compare RT with a Best-Single-Model Policy (BSMP), which simply selects the response from the single model with the highest confidence weight rather than synthesizing across multiple players. Conceptually, BSMP resembles a committee that crowns one “top performer” and ignores all other voices. The results in Figure 4 make clear why this approach falls short: across both MultiTask and SingleTask settings, RT consistently outperforms BSMP. In MultiTask, the advantages are particularly evident in reasoning-heavy domains such as mathematics and mechanics, where drawing on complementary perspectives yields large gains. In SingleTask proposal writing, consensus fusion not only raises average scores across all rubric dimensions but also reduces variance, producing outputs that are both stronger and more consistent. These findings highlight that scientific progress rarely emerges from a single dominant voice; it is the synthesis of multiple viewpoints that produces reliable reasoning and coherent narrative.
Rounds Ablation.
We next examine the impact of the number of training rounds used to construct the confidence-weight table. Here, the table plays the role of institutional memory—accumulating evidence about which players are dependable under different conditions. As shown in Figure 5, the effect of this memory differs by task type. In MultiTask, where benchmarks span diverse scientific domains, performance steadily improves as more rounds are incorporated and stabilizes only after hundreds of evaluations. By contrast, SingleTask writing converges much more quickly: fewer than 100 rounds already suffice to surpass baselines, with further rounds offering only modest additional benefit. This divergence reflects a natural distinction: heterogeneous domains require extensive calibration to build trust across disciplines, whereas rubric-based writing tasks benefit from faster stabilization since evaluative criteria are more homogeneous.
Summary.
These ablations underscore two core insights: First, weighted consensus is fundamentally more powerful than relying on a single “best” model, as robust science arises from collective deliberation rather than solitary voices. Second, the effectiveness of the confidence-weight table depends on task structure: structured MultiTask settings demand long-run calibration, while open-ended SingleTask writing achieves stability with much less historical data. Together, these findings reinforce that consensus-driven fusion is not only effective but also principled—mirroring how real scientific communities build reliable knowledge through collective judgment.

### v2

@@ 3.1 Experimental Setup
| ScienceNarrative |
GPT-4o
70.995.01 | 74.734.68 | 82.753.76 | 75.495.23 | 74.815.03 | 79.634.31 | 72.155.19 | 70.535.25 | 79.644.19 |
GPT-o3
84.073.98 | 84.253.82 | 86.463.42 | 85.923.95 | 86.703.56 | 85.843.56 | 83.904.15 | 81.524.25 | 84.863.69 | |
GPT-4.1
76.964.62 | 81.374.16 | 86.193.51 | 81.654.67 | 83.194.19 | 84.753.85 | 78.814.74 | 77.914.67 | 83.903.88 | |
GPT-4.1n
77.394.55 | 80.944.15 | 85.593.53 | 80.484.69 | 82.304.19 | 83.973.85 | 77.974.74 | 77.224.65 | 82.943.90 | |
GPT-o3-m
82.654.16 | 84.783.83 | 87.263.36 | 83.154.35 | 86.863.53 | 86.443.53 | 81.504.40 | 81.824.23 | 85.153.67 | |
GPT-4o-m
72.784.94 | 74.644.70 | 82.193.80 | 75.445.16 | 76.184.85 | 79.814.23 | 72.785.18 | 70.695.24 | 78.824.25 | |
Grok-3
78.505.10 | 79.104.50 | 85.503.64 | 80.964.75 | 81.264.47 | 83.583.99 | 78.814.78 | 76.344.93 | 83.044.01 | |
Claude-3.7S
82.474.22 | 84.733.92 | 88.173.28 | 85.524.20 | 86.653.79 | 86.893.57 | 84.534.22 | 82.134.38 | 86.413.59 | |
DeepSeek-chat
84.054.20 | 83.744.12 | 88.283.19 | 84.214.31 | 84.584.13 | 86.083.58 | 83.684.23 | 81.564.48 | 86.503.46 | |
| Ours(RP) | 90.253.95 | 89.953.78 | 91.043.09 | 90.393.91 | 91.283.35 | 91.183.20 | 90.153.95 | 89.863.95 | 90.993.26 | |
ScienceEval: Agents are simultaneously queried with nine domain-specific questions spanning geoscience (Deng et al., 2024), biology (Jin et al., 2019), mathematics (Cobbe et al., 2021), and multi-faceted physical science(Appendix §C), containing six sub-categories. The tasks include multiple-choice, fill-in-the-blank, and short-answer formats. This setting resembles a comprehensive written examination covering diverse scientific disciplines, where success requires broad and balanced reasoning capabilities rather than specialization in a single domain.
ScienceNarrative: Agents are required to generate structured scientific proposals following a standardized three-paragraph template consisting of Background, Methodology, and Impact (§C). This regime emphasizes the narrative dimension of scientific reasoning: beyond producing isolated facts, agents must organize ideas into a coherent, logically consistent, and persuasive scientific narrative.
Baselines. We evaluate Roundtable Policy against both single-agent and multi-agent baselines. As single-agent baselines, we adopt a diverse set of state-of-the-art LLMs following the notation in Table 4. For multi-agent collaboration, we consider representative frameworks: Majority Vote (Wang et al., 2023) aggregates agents by simple voting and serves as a canonical non-weighted baseline. Weighted Vote (Taubenfeld et al., 2025) extends majority voting by assigning confidence-based weights to individual agents. Debate (Du et al., 2024) enables agents to iteratively revise their answers through argument exchange. Debate w Judge (Liang et al., 2024) introduces an explicit adjudicator to select the final answer after debate, while Debate w Confidence (Chen et al., 2024) further incorporates confidence signals to guide resolution. Additionally, Mixture-of-Agents (Wang et al., 2025) constructs an agentic network to combine multiple agent outputs adaptively.
3.2 Performance Improvement (RQ1)
To evaluate the efficacy, we compare Roundtable Policy (RP) against single-LLM baselines across both the ScienceEval and ScienceNarrative. Table 1 summarizes overall performance, while Figures 13 and 15 visualize fine-grained percentage improvements. Additional per-player and per-task analyses are provided in Appendix §D.1 and §D.2. Based on these results, we draw the observations:
Obs 1: RP consistently improves performance across heterogeneous scientific tasks, with particularly strong gains on reasoning-intensive problems, while maintaining stable and reliable outcomes. In the ScienceEval, RP achieves a 13.01% average improvement over single-model baselines. Performance gains are broadly positive on Tasks 1–8 and are especially pronounced on reasoning-heavy subtasks (multi-faceted physical reasoning). Beyond mean accuracy, RP also reduces predictive uncertainty; for example, on Task 7 (Math), RP narrows the confidence interval to , compared with – for single-model baselines. In contrast, Task 9 (Biology) exhibits marginal or negative differences, likely due to high intrinsic uncertainty and substantial disagreement among individual models, which limits the effectiveness of consensus-based aggregation. These results indicate that RP is most effective when model errors are diverse yet structured.
Obs 2: RP yields robust and uniform improvements in structured scientific writing, substantially elevating weaker drafts without degrading strong ones. In the ScienceNarrative, RP achieves superior performance across all nine rubric dimensions, with an 11.04% average improvement. Gains concentrate in the Impact section—particularly Scientific Rigor—while overall scores remain tightly clustered around 90–91% with narrow confidence intervals. Per-player analyses reveal right-skewed improvement distributions, indicating that RP consistently lifts weaker proposals while preserving the quality of already strong drafts.
3.3 Stability and Robustness (RQ2)
Setting
Model
Avg. Gain (%)
Std (%)
p-value
Significance
ScienceEval
-4o
6.476865
77.280369
-2
5.254293
78.231667
-3.5H
17.251754
101.839946
-4.5p
18.552776
94.463211
-Chat
16.048551
96.935292
-3.7S
13.732040
121.321577
-3.3
24.988341
123.678543
-plus
12.586367
91.751670
-2.0
46.474125
145.071773
ScienceNarrative
-4o
27.087781
40.303119
-o3
13.316247
148.785675
-4.1
15.107889
73.600957
-4.1-nano
15.805949
28.188351
-o3-mini
9.687830
21.760829
-4o-mini
27.206693
42.707340
-3
16.485277
29.024328
-3.7S
7.761157
16.993190
-Chat
7.165863
7.474061
We examine whether the performance gains are statistically reliable and robust under long-run interactions and diverse design choices. Specifically, we conduct Wilcoxon signed-rank tests to assess statistical significance and perform systematic ablations to evaluate structural robustness (Figure 5).
Obs 3: RP provides statistically reliable and long-run stable performance improvements. As shown in Table 2, RP significantly outperforms all agents at , with -values typically below . Despite substantial per-round variance, the aggregated gains consistently converge to stable improvements over of thousands of evaluations. This phenomenon indicates that RP’s advantages are not driven by isolated cases or short-lived fluctuations, but reflect persistent distribution-level improvements. Fine-grained Wilcoxon results are reported in Tables 11, 12, 9 and 10.
Obs 4: RP’s stability is structurally grounded and manifests consistently across multiple design parameters. Figure 5 shows that RP remains robust under systematic ablations of various components. Scaling the number of player agents leads to steady performance improvements that gradually saturate. In practice, the number of players is constrained by the maximum context window of the aggregator, as all proposals are concatenated for aggregation; we therefore evaluate scalability within this realistic operating regime. Second, ablating the aggregator from strong to weak results in bounded performance degradation rather than drastic collapse, indicating that RP does not rely on a single highly capable aggregator. Additionally, increasing the number of pretraining rounds111historical rounds to distill the confidence-weight table yields consistent gains, indicating that more accurate encoding of historical reliability leads to better downstream performance. Finally, incorporating uncertainty-aware scoring outperforms vanilla score aggregation, reflecting that RP’s reasoning improvements are not driven by fragile or ad hoc scoring heuristics.
3.4 Multi-Agent Baselines (RQ3)
| Setting | Method | Multi-Faceted Physical Reasoning | Other Domains | |||||||
| Mech. 1 | Mech. 2 | EE | Opt. | Thermo. | Semi. | Math | Geo. | Bio. | ||
| ScienceEval | Majority Vote (Wang et al., 2023) | 82.484.51 | 74.025.63 | 85.014.25 | 73.655.48 | 81.014.56 | 81.684.77 | 85.932.12 | 79.032.27 | 50.886.54 |
| Weighted Vote (Taubenfeld et al., 2025) | 81.564.58 | 74.465.54 | 84.244.29 | 73.135.45 | 81.764.49 | 81.164.77 | 83.032.27 | 76.562.31 | 49.286.48 | |
| Debate (Du et al., 2024) | 85.684.12 | 80.344.28 | 88.014.02 | 77.253.90 | 84.973.30 | 89.174.09 | 92.231.05 | 84.051.90 | 51.686.25 | |
| Debate w Judge (Liang et al., 2024) | 86.934.28 | 80.364.20 | 88.424.10 | 76.254.05 | 85.263.77 | 89.434.25 | 92.041.25 | 83.651.97 | 51.236.07 | |
| Debate w Confidence (Chen et al., 2024) | 86.243.50 | 79.613.69 | 88.153.55 | 79.393.73 | 86.273.02 | 89.953.45 | 93.181.55 | 83.921.30 | 49.174.75 | |
| Mixture-of-Agents (Wang et al., 2025) | 87.533.75 | 81.023.74 | 89.743.47 | 80.683.23 | 87.293.05 | 91.253.59 | 93.051.42 | 85.281.25 | 53.505.31 | |
| Roundtable Policy | 87.644.26 | 81.505.36 | 89.994.00 | 81.294.96 | 86.554.09 | 91.824.34 | 94.801.01 | 85.331.63 | 47.557.02 | |
| Setting | Model | Background | Methodology | Impact | ||||||
| ScienceNarrative | Majority Vote (Wang et al., 2023) | 81.374.30 | 82.994.07 | 86.963.41 | 83.464.37 | 84.643.98 | 85.433.70 | 81.934.41 | 80.114.51 | 84.903.73 |
| Weighted Vote (Taubenfeld et al., 2025) | 81.474.30 | 83.224.05 | 87.263.36 | 83.744.36 | 84.903.97 | 85.743.68 | 82.374.38 | 80.354.51 | 85.223.70 | |
| Debate (Du et al., 2024) | 83.164.88 | 84.995.25 | 88.253.64 | 86.145.25 | 87.203.50 | 86.923.25 | 85.114.02 | 83.654.00 | 85.963.75 | |
| Debate w Judge (Liang et al., 2024) | 86.753.75 | 85.504.25 | 86.044.17 | 87.053.50 | 88.963.25 | 88.813.10 | 87.254.06 | 83.503.85 | 85.413.71 | |
| Debate w Confidence (Chen et al., 2024) | 86.473.20 | 89.253.34 | 90.022.95 | 86.993.13 | 90.253.25 | 90.702.75 | 87.083.50 | 88.633.31 | 89.822.95 | |
| Mixture-of-Agents (Wang et al., 2025) | 89.983.90 | 86.174.47 | 89.503.68 | 90.253.65 | 88.923.25 | 90.153.57 | 89.754.25 | 88.603.93 | 89.223.10 | |
| Roundtable Policy | 90.253.95 | 89.953.78 | 91.043.09 | 90.393.91 | 91.283.35 | 91.183.20 | 90.153.95 | 89.863.95 | 90.993.26 | |
Furthermore, we compare Roundtable Policy with representative multi-agent frameworks (implementation details at Appendix §E), including voting-based aggregation (Wang et al., 2023; Taubenfeld et al., 2025), debate-style interaction (Du et al., 2024; Liang et al., 2024; Chen et al., 2024), and agentic-network-style aggregation (Wang et al., 2025). Beyond raw performance, we analyze their behaviors through the lens of how agent influence is assigned, updated, and stabilized during aggregation.
Obs 5: RP achieves consistently strong and balanced performance in complex reasoning environments. In the ScienceEval, RP attains the best or second-best results on 8 out of 9 task dimensions, with particularly strong performance on reasoning-intensive domains such as multi-faceted physical reasoning and Math. This behavior contrasts with voting-based methods, whose uniform and memoryless aggregation is prone to majority bias when partial but confident opinions dominate. While interaction-heavy baselines such as Debate and Debate w Confidence can achieve competitive or even superior performance on individual tasks, their gains are highly task-dependent and exhibit noticeable instability, reflecting sensitivity to transient interaction dynamics. The agentic-network-style method (Mixture-of-Agents) also demonstrates strong performance, which we attribute to the rich architectural expressivity: layered agents integrate diverse perspectives, analogous to a highly expressive super-neuron network. However, this expressivity comes with structural redundancy, and the aggregation mechanism remains largely static, limiting its ability to adaptively recalibrate agent influence under diverse environments.
Obs 6: RP demonstrates superior generality in complex scientific writing. In the ScienceNarrative, RP achieves best performance across all nine rubrics. Debate-based methods excel in dimensions emphasizing argumentative clarity or coherence, but their performance degrades when tasks require creativity. Debate with Confidence partially mitigates this issue by incorporating confidence cues during interaction; however, such cues remain interaction-conditioned and often lead to over-neutralization. As shown in Table 3, competing methods exhibit strengths in specific categories, whereas RP consistently maintains balanced performance across all rubrics.
3.5 Bias and Consistency (RQ4)
We analyze grader bias and inter-grader consistency using (i) rank distributions assigned to RP and (ii) pairwise agreement measured by Kendall’s Tau (Appendix §A.3). Aggregate results are shown in Figure 7, with detailed breakdowns in Appendix §D.3.
Obs 7: AI graders are consistent on structured tasks and more variable on open-ended evaluations, while committee aggregation mitigates individual bias. In ScienceEval, graders exhibit concentrated rank distributions and high Kendall’s Tau values, indicating strong agreement on objective tasks. In ScienceNarrative, inter-grader agreement decreases, particularly for subjective dimensions such as creativity and logical coherence. Nevertheless, aggregating across multiple graders substantially reduces individual bias, yielding balanced assessments.
Summary. By decoupling influence assignment from transient interaction dynamics and aligning long-term fine-grained measurement, RP enables stable context-aware aggregation. In addition to increasingly complex interaction protocols, our observations find that multi-agent systems also benefit from mechanisms that explicitly model, update, and stabilize agent reliability across downstream settings. Future inference-time reasoning frameworks for multi-agent systems may increasingly view collaboration as a problem of reliability modeling and consensus formation, rather than relying solely on more elaborate interaction dynamics.